OpenStudy (anonymous):

Really need help with this problem . Solve the following problem using substitution or elimination: To skate at Roller Heaven, each person must pay a monthly membership fee (m) and a fee for each session (f). In January, Art attended 5 sessions and paid $25 in all. In July, he attended 10 sessions, and he paid $40 in all. How much is the membership fee and how much are the session? Be sure to show all work.

OpenStudy (anonymous):

They helped you by defining the variables in the problem, so set up the system: \[m+5f=25\]\[m+10f=40\]then solve the system.

OpenStudy (anonymous):

Take the second equation, ans subtract the first to get \[5f=15\rightarrow f=3\rightarrow m=10\]so our solution is\[(m,f)=(10,3)\]

hero (hero):

m = 25 - 5f m = 40 - 10f m = m 25 - 5f = 40 - 10f 10f - 5f = 40 - 25 5f = 15 f = 15/5 f = 3 m = 10

OpenStudy (anonymous):

Thanks guys !!!